One mole of an ideal gas expands reversibly from a volume of 5 dm^3 and temperature of 25 °C to a volume of 10 dm^3 and temperature of -23 °C. Calculate the entropy change accompanying the process. (Assume C v= 1.5R, R= Universal gas constant).


1
Expert's answer
2021-02-15T03:22:44-0500

"\Delta"S = "\Delta"ST + "\Delta"SP = nCpln(T2 / T1) + nRln(V2 / V1) = n(Cv + R)ln(T2 / T1) + nRln(V2 / V1) = 9.814ln(250 / 298) + 8.314ln(10 / 5) = 4.04 (J / K).

Answer: "\Delta"S = 4.04 (J / K)